Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-11-21

Observers: K. Takáts, Ósmar Rodríguez
Support team: Lluís Galbany
TAT team: Joe Anderson, Kate Maguire

Daytime calibrations:
EFOSC2:
PESSTO_bias
PESSTO_calib

The night is not photometric

Note: telescope was closed until 01:20 UT because of the weather. We couldn't take evening skyflats

Weather conditions:
Start: Clouds, high humidity, moderate wind, pointing restriction
Middle: Thin clouds, moderate wind, pointing restriction
End: Thin clouds, moderate wind, pointing restriction

01:26 Image analysis
01:25 Spectroscopic standard vma2, seeing ~1.3"-1.7"
02:00 SN2016iae follow-up, gr11+gr16 with 1.5" slit,  seeing ~1.5" (gr11), 1.2"-1.6"(gr16)
03:15 OGLE16euo follow-up, gr16 with 1.5" slit, seeing ~1.7"-2.4"
04:15 OGLE16euo follow-up, gr16 with 1.5" slit, seeing ~1.9"-2.5"
05:15 OGLE16fnh classification, seeing ~1.8"-2.0"
05:55 OGLE-2015-SN-035 follow-up, BVRi, seeing ~1.9" (B), ~1.8"(V), ~1.5"(R), ~1.6"(i)
08:00 Spectroscopic standard LTT3864, only with the 1.5" slit, seeing ~1.6"
08:25 End of the night
Comments